OpenStudy (anonymous):

what is the simplified form of x/x^2-y^2 plus y/x^2-y^2

ganeshie8 (ganeshie8):

when denominators are same, you can add numerators

ganeshie8 (ganeshie8):

\(\large \frac{x}{x^2-y^2} + \frac{y}{x^2-y^2} \)

OpenStudy (anonymous):

i thought it was x+y/x^2-y^2. but its not any of my answer choices

ganeshie8 (ganeshie8):

thats right ! we just need to simplify further

ganeshie8 (ganeshie8):

\(\large \frac{x+y}{x^2-y^2} \)

ganeshie8 (ganeshie8):

we can factor denominator !!

OpenStudy (anonymous):

soo will it be 1/x-y ?

ganeshie8 (ganeshie8):

Yes !! very good work :)

OpenStudy (anonymous):

omg thank you verryy muchh (:

ganeshie8 (ganeshie8):

np :)